evaluate :
(-4+6) -3 (-2)

Found 2 solutions by rfer, math-vortex:
Answer by rfer(12671)
Answer by math-vortex(472)

Evaluate:
 (-4+6) -3(-2) =

First, we want to clear the parentheses. (-4+6) is 2.
2 -3(-2) =

2 -3(-2) means -3 times -2, or +6. (Keep the signs with the numbers.
2+6 = 8

Your expression simplifies to 8.

Hope this helps!

Ms.Figge
math.in.the.vortex@gmail.com